OpenStudy (anonymous):

Mike is saving to buy a car. He already has $500 in the bank, and he plans to save $250 each month. This relationship is represented by the equation y=250x+500, where x is the number of months and y is the total money saved. 1. Identify the slope. 2. Explain what the slope means in this problem. 3. Identify the y-intercept. 4. Explain what the y-intercept means in this problem.

OpenStudy (anonymous):

First remeber that the standard from of a line is Y = Mx + C

OpenStudy (anonymous):

1. M is the slope so that is 250

OpenStudy (anonymous):

2. slope is a rate of change - so in this equation - he is planning to save at a rate of 250 per month

OpenStudy (anonymous):

3. Y intercept are where x = 0 y = 250(0) +500 y= 500 intercept is (0, 500)

OpenStudy (anonymous):

4. y intecerpt is where the lines crosses the y axis.

OpenStudy (anonymous):

All about y = mx + b Total payment = monthly payment + deposit